# KOALA++: Efficient Kalman-Based Optimization of Neural Networks with Gradient-Covariance Products

### v1

@@ C.1 Effect of Batch Size on KOALA++ Performance
We evaluate the impact of batch size on the performance of KOALA++ using ResNet50 on the CIFAR100 dataset. As shown in Table 6, increasing the batch size leads to a noticeable degradation in both Top-1 and Top-5 accuracy. The results suggest that KOALA++ is more effective in smaller-batch training regimes, possibly due to better gradient variance estimation and more frequent updates.
@@ C.2 Different Schedulers
C.3 Different Initializations of KOALA++
@@
Table 8: Effect of varying $\sigma$ on KOALA++ performance. Dataset: CIFAR-100, Model: ResNet-50, Epochs: 200, $q=0.2$, lr$=2.0$
@@
We observe that increasing $\sigma$ leads to improved performance, suggesting that initializing $v_{1}$ with stronger magnitude provides better early-stage directionality, which helps escape suboptimal regions in the optimization landscape. The result is shown in Figure 5
@@
Appendix D Experiments Details
D.1 Hardware
@@
D.2 Image Classification
D.2.1 HP Settings for CIFAR10/100
We evaluate KOALA++ on both CIFAR-10 and CIFAR-100 using a variety of architectures, including CNN-based models (e.g., ResNet [8]) and vision transformers (e.g., Swin-Tiny [16]). For CIFAR-10, we initialize the process noise $\sigma$ and uncertainty level $q$ to 0.1. The initial learning rate is set to 1.0. In contrast, for CIFAR-100, given the larger number of classes and relatively richer sample distribution, we adopt a more aggressive hyperparameter configuration: both $\sigma$ and $q$ are set to 0.2, and the initial learning rate is increased to 2.0. For both datasets, we apply a weight decay of $5\times 10^{-4}$ to all ResNet and other convolutional models.
When applying KOALA++ to Swin-Tiny, we perform a coarse grid search over the weight decay parameter, evaluating the values $\{1\times 10^{-1},1\times 10^{-2},5\times 10^{-3},1\times 10^{-4},5\times 10^{-4}\}$. Interestingly, we find that KOALA++ tends to perform best with a smaller weight decay value ($1\times 10^{-4}$), whereas other optimizers usually favor a higher decay value (e.g., $1\times 10^{-2}$).
We hypothesize that this is due to KOALA++’s implicit regularization behavior stemming from its dynamic covariance updates, which already exert a stabilizing effect on parameter updates. Therefore, combining KOALA++ with a strong explicit regularizer such as large weight decay may lead to underfitting, particularly in transformer-based models where overparameterization is more common. Hence, a smaller weight decay appears more suitable when using KOALA++ in this context.
D.2.2 More Results for CIFAR10/100
@@ D.2.3 Experiments for ImageNet32
D.3 Language Modeling

### v2

@@ D.1 Effect of Batch Size on KOALA++ Performance
We evaluated the impact of batch size on the performance of KOALA++ using ResNet50 on the CIFAR100 dataset. As shown in Table 6, increasing the batch size leads to a noticeable degradation in both Top-1 and Top-5 accuracy. The results suggest that KOALA++ is more effective in smaller-batch training regimes, possibly due to better gradient variance estimation and more frequent updates.
@@ D.2 Different Schedulers
D.3 Different Initializations of KOALA++
@@
Table 8: Effect of varying $\sigma$ on KOALA++ performance. Dataset: CIFAR-100, Model: ResNet-50, Epochs: 200, $q=0.2$, lr=2.0=2.0
@@
We observe that increasing $\sigma$ leads to improved performance, suggesting that initializing $v_{1}$ with stronger magnitude provides better early-stage directionality, which helps escape suboptimal regions in the optimization landscape. The results are shown in Figure 7.
@@
Appendix E Experiments Details
E.1 Hardware
@@
E.2 Image Classification
E.2.1 HP Settings for CIFAR-10/100 and Hyperparameter Sensitivity
We evaluate KOALA++ on both CIFAR-10 and CIFAR-100 using a variety of architectures, including CNN-based models (e.g., ResNet [9]) and vision transformers (e.g., Swin-Tiny [18]). Across all experiments, we used the same hyperparameter settings for CNN-based models to highlight KOALA++’s low-tuning usability, as it achieves competitive or better results than baselines without dataset-specific tuning. When moving to Transformer-based architectures, we conducted a small grid search following the procedure described in the Supplementary Material.
Below we provide further clarification on how the key hyperparameters are selected, consistent with the analysis included in the Supplementary Material to support practical deployment of KOALA++:
  • Measurement noise $R$: estimated online as an exponential moving average of the mini-batch loss variance, using a smoothing factor $\alpha=0.9$. This removes $R$ from the set of hyperparameters requiring manual tuning.
  • Initial variance $\sigma_{0}$: controls only the scale of the initial directional state $v_{1}=\sigma_{0}H_{1}$. Ablation studies indicate that KOALA++’s convergence is robust to this value, so we fix $\sigma_{0}$ to match the process noise variance $Q$ for simplicity and symmetry.
  • Process noise $Q$: the main hyperparameter that requires tuning besides the learning rate. Empirically, we found that $Q\in[0.1,0.4]$ performs well across datasets.
For CIFAR-10, we initialize both $\sigma_{0}$ and $Q$ to $0.1$, with an initial learning rate of $1.0$. For CIFAR-100, which has more classes and a richer data distribution, we adopt slightly larger values $\sigma_{0}=Q=0.2$ and increase the initial learning rate to $2.0$. A weight decay of $5\times 10^{-4}$ is applied to all ResNet and other CNN models.
When applying KOALA++ to Swin-Tiny, we perform a coarse grid search over the weight decay parameter, evaluating values $\{1\times 10^{-1},1\times 10^{-2},5\times 10^{-3},1\times 10^{-4},5\times 10^{-4}\}$. Interestingly, KOALA++ performs best with a smaller weight decay ($1\times 10^{-4}$), whereas other optimizers typically favor higher decay values (e.g., $1\times 10^{-2}$). We hypothesize that this stems from KOALA++’s implicit regularization through dynamic covariance updates, which already stabilize parameter updates. Thus, combining KOALA++ with a large explicit regularizer (e.g., strong weight decay) can lead to underfitting—especially in transformer-based models—making a smaller weight decay more suitable in such cases.
E.2.2 More Results for CIFAR10/100
@@ E.2.3 Experiments for ImageNet32
E.3 Efficiency Analysis
We provide additional results on the computational efficiency of KOALA++. Besides accuracy, it is important that an optimizer remains efficient in terms of runtime and memory, particularly when scaling to larger models and batch sizes.
ResNet-50 and Swin-Tiny (batch size 256).
Table 10 reports average training time per epoch, asymptotic FLOPs, and peak memory usage on ResNet-50 and Swin-Tiny. KOALA++ achieves runtime and memory comparable to Adam(W), while being substantially more efficient than Shampoo and AdaHessian. For K-FAC, we note that the more precise per-iteration complexity is $O(\ell d^{2}m)$ [^8] Here $\ell$ is the number of layers, $d$ the typical hidden dimension, and $m$ the batch size., although it is commonly reported as $O(n^{2})$ in the literature.
Table 10: Efficiency profiling on ResNet-50 and Swin-Tiny (batch size $256$). We report average training time per epoch, asymptotic FLOPs, and peak memory usage.
ResNet-50 (Batch size 256)
Optimizer Time/epoch (s) FLOPs (class) Peak Mem (GB)
Adam 11.08 $O(n)$ 10.51
AdaFisher 11.64 $O(n)$ 11.91
AdaHessian [^9] AdaHessian uses Hutchinson’s trick for Hessian diagonal estimation, which requires additional Hessian-vector products, leading to $O(n^{2})$ complexity. 63.43 $O(n^{2})$ 22.79
K-FAC [^10] More precisely $O(\ell d^{2}m)$, where $\ell$ is the number of layers, $d$ the hidden dimension, and $m$ the batch size. 17.16 $O(n^{2})$ 12.34
Shampoo 134.63 $O(n^{1.5})$ 10.41
KOALA++ (ours) 13.23 $O(n)$ 10.27
Swin-Tiny (Batch size 256)
Adam(W) 16.35 $O(n)$ 3.46
AdaFisher 12.00 $O(n)$ 4.38
AdaHessian 42.00 $O(n^{2})$ 5.97
KOALA++ (ours) 16.10 $O(n)$ 4.50
ResNet-101 and Swin-Base (batch size 1024).
To further test scalability, we report results on ResNet-101 and Swin Transformer Base with batch size $1024$ in Table 11. KOALA++ remains competitive with Adam(W), while scaling better than AdaHessian and Shampoo in both runtime and memory. This confirms that KOALA++ maintains near-first-order efficiency even at larger scales.
Table 11: Efficiency profiling on ResNet-101 and Swin Transformer Base with batch size $1024$.
ResNet-101 (Batch size 1024)
Optimizer Time/epoch (s) FLOPs (class) Peak Mem (GB) Adam 16.79 $O(n)$ 48.16 AdaFisher 16.62 $O(n)$ 48.11 K-FAC 23.79 $O(n^{2})$ 52.09 Shampoo 64.29 $O(n^{1.5})$ 48.90 KOALA++ (ours) 17.61 $O(n)$ 48.25
Swin Transformer Base (Batch size 1024)
Optimizer Time/epoch (s) FLOPs (class) Peak Mem (GB) Adam(W) 19.07 $O(n)$ 23.23 AdaFisherW 18.95 $O(n)$ 22.28 AdaHessian 57.72 $O(n^{2})$ 64.08 KOALA++ (ours) 19.52 $O(n)$ 22.51
Comparison with SGD across batch sizes.
We also compare throughput with SGD on CIFAR-10 using ResNet-50 under varying batch sizes (Table 12). KOALA++ incurs a small constant overhead per iteration due to $O(n)$ covariance–vector products, but this cost is independent of batch size. As the batch size grows, forward/backward cost dominates, and the relative overhead of KOALA++ shrinks, approaching parity with SGD at batch size $1024$.
Table 12: Throughput comparison of KOALA++ vs SGD across different batch sizes on CIFAR-10 with ResNet-50.
Batch Size SGD (s/epoch) KOALA++ (s/epoch) Ratio (KOALA++ / SGD) Overhead
[…]
This trend aligns with our design intuition: within each batch, the effective covariance dimension becomes smaller as stochastic noise is averaged out, so the extra operations of KOALA++ contribute less to the total runtime. Consequently, KOALA++ scales gracefully with batch size and approaches the efficiency of SGD in large-batch regimes.
E.4 Language Modeling